A block hey of mass 2 kilogram rest on another block B of mass 8 kilogram which rest on original floor the coefficient of friction between a and b is 0.2 while that between be and floor is 0.5 when a horizontal force of 25 newton is applied on the block B the force of friction between a and b is

Grade:11

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
Block B together with block A is lying on rough surface with coefficient of friction=0.5.
So, it will be able to accelerate if force applied is greater than (coefficient of friction)*(mA+mB)*g=0.5*10*10=50 N.
But force applied is only 25 N which is less than 50 N.
So there will be no acceleration and frictional force between A and B wont arise.
So frictional force= 0 [Ans]
Regards,
Soumya Ranjan Mohanty.